OpenStudy (anonymous):

In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?

OpenStudy (anonymous):

A=90 B=30 C=60

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

write an equation a=3b c=90-b 3b+b+90-b=180 3b+90=180 b=30 a=3b=90 b+c=90==>c=60

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you